# Smart Exploration in Reinforcement Learning Using Bounded Uncertainty Models

### v2

@@ III-B1 Choice of Weighting $\beta$
$\beta(x,u)=\text{Prob}[I(x,u)>0]=\frac{\max\left(0,\bar{Q}(x,u)-\underaccent{\bar}{V}(x)\right)}{\bar{Q}(x,u)-\underaccent{\bar}{Q}(x,u)}.$ (13)
@@
$\beta(x,u)=\mathbb{E}[I(x,u)]=\frac{\left(\max\left(0,\bar{Q}(x,u)-\underaccent{\bar}{V}(x)\right)\right)^{2}}{2\left(\bar{Q}(x,u)-\underaccent{\bar}{Q}(x,u)\right)}.$ (14)
@@ III-C Regularization with Observed Data
We propose that by regularizing the optimization over the model set using observed data $\mathcal{D}$, the bounds converge to the optimal Q-function under mild conditions. We can re-define the Q-bound Bellman updates (4) as
@@
The variable $\lambda$ determines the trade-off between the model set and the observed data. It should increase with the size of $\mathcal{D}$ to ensure that the selected model from the set is biased towards the observed data, such that
@@
where $T(x,u,x^{\prime})$ is the number of observed transitions from state $x$ under action $u$ to state $x^{\prime}$ in the dataset $\mathcal{D}$, and $T(x,u):=\sum_{x^{\prime}}T(x,u,x^{\prime})$. In practice, $\lambda$ can be chosen as any increasing function of the visit count $T(x,u)$ that satisfies (17).
Since the rewards are deterministic, whenever a reward $r_{k}$ is observed at state-action pair $(x_{k},u_{k})$, the set of reward functions $\mathcal{G}$ can be constrained to include only elements that correctly assign this reward at that state-action pair, i.e.,
$\mathcal{G}=\{g\in\mathcal{G}\mid g(x_{k},u_{k})=r_{k}\}.$ (18)
Together with the regularized model set optimization, we obtain a convergence result which is formalized next.
@@
Assuming finite state-action spaces, let $\lambda:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{\geq 0}$ be a regularization parameter satisfying (17), and let $d(P,P_{E})$ be a distance metric such that $d(P,P_{E})\geq 0$ and $d(P,P_{E})=0$ if and only if $P=P_{E}$. Let $\mathcal{P}$ and $\mathcal{G}$ be the set of transition kernels and the set of reward functions, respectively. Assume that the reward function is deterministic. If every state-action pair is visited infinitely often, then,
@@
Since every state-action pair is visited infinitely often, $\lambda(x,u)\to\infty$ by (17). The regularization term then forces the optimization to select $P=P_{E}$, as the regularization dominates the objective. With deterministic rewards, $\mathcal{G}$ reduces to the true reward function after one visit per state-action pair through (18). By the law of large numbers, $P_{E}$ converges to the true transition kernel $P$. As the Bellman operator is a contraction (see [13]), it follows that both the lower and upper Q-function bounds converge to $Q^{*}$. ∎
@@
Algorithm 1 BUMEX: Q-learning with exploring policy and regularized model set optimization
@@
   Clip the Q-function $Q$ to the bounds $\underaccent{\bar}{Q},\bar{Q}$.
@@
The regularized model set optimization is incorporated into our proposed method, which is summarized in Algorithm 1, where the data-regularized Q-bound iterations are performed every $L$ episodes. To improve learning efficiency, we employ an $\epsilon$-greedy type strategy, balancing exploration/exploitation. Additionally, we clip the Q-function to the bounds whenever they are updated. By combining our previous results, we obtain the following convergence guarantee for the exploring policy.
@@
Assume finite state and action spaces. Let $\pi^{\phi}$ be the exploring policy defined in (10) for which we additionally assume that $\zeta$ is positive. Let $\lambda$ satisfy (17). Assume that $d(P,P_{E})$ is a distance metric with $d(P,P_{E})\geq 0$ and $d(P,P_{E})=0$ if and only if $P=P_{E}$. Assume deterministic rewards, and that every state is visited infinitely often. Then, under Algorithm 1, the Q-bounds converge to the optimal Q-function and the exploring policy $\pi^{\phi}$ converges to the optimal policy with probability 1 as $k\to\infty$.
@@
With $\zeta>0$, the exploring policy assigns positive weight to all actions except those that are certain and guaranteed suboptimal. For uncertain state-action pairs, the policy ensures infinite exploration, causing $\lambda(x,u)\to\infty$ and forcing convergence to the empirical kernel. By Proposition 5, the Q-bounds converge to $Q^{*}$. The exploring policy then selects only guaranteed optimal actions, converging to the optimal policy. ∎
IV PRACTICAL ALGORITHM: BUMEX
@@
for known, well-defined lower and upper bounds $\underaccent{\bar}{M},\bar{M}\in[0,1]^{|\mathcal{X}|\times|\mathcal{U}|\times|\mathcal{X}|}$. The reward function set that the true reward function $g$ belongs to is defined as $\mathcal{G}=\left\{g~\big|~\underaccent{\bar}{g}\leq g\leq\bar{g},\quad\forall~x,u\right\}$. Thus, we assume that the true transition kernel and reward function are bounded by known upper and lower bounds.
The distance metric $d$ is chosen to be the Kullback-Leibler (KL) divergence between the transition kernel and the observed data, given by
@@
where $M_{E}$ is the empirical transition probability tensor from the observed data $\mathcal{D}$ through $M_{E}(x,u,x^{\prime})=\frac{T(x,u,x^{\prime})}{T(x,u)}$. We choose the KL-divergence since it satisfies the nonnegativity and zero-distance properties, is convex, and is a natural choice for comparing probability distributions.
@@
We propose to select $\lambda=0$ in the absence of data ($T(x,u)=0$) and otherwise as a function of system properties and the number of observed transitions from every state-action pair, such that
@@
The algorithm presented in this section is a straightforward extension of the tabular Q-learning algorithm, where we can use the Q-function bounds to guide the agent’s exploration. The Q-bound updates are computationally efficient: without regularization, they reduce to simple min/max operations over the model set bounds, while with regularization, the optimization remains convex and tractable. The computational complexity scales similarly to other finite MDP methods like UCRL2 or PSRL, with runtime dominated by the state-action space size rather than the model set optimization. Note that all the convergence properties of Q-learning discussed in the previous sections are retained. In some cases, we can even guarantee finite-time convergence of the exploring policy to the optimal policy, as shown in the next section.
@@ IV-A Finite-Time Convergence
Furthermore, $\lambda(x,u)$ is set to rely entirely on the data for the deterministic case, that is, $\lambda(x,u)=\infty$ if 1) holds and $T(x,u)>0$.
@@
Let $\pi^{\phi}$ be the exploring policy defined in (10) and let $Q$ be the Q-function from Algorithm 1, using model set $\hat{\mathcal{M}}$ and distance metric $d$ with $d(P,P_{E})\geq 0$ and $d(P,P_{E})=0$ if and only if $P=P_{E}$. Assume deterministic rewards. Assume that every state is visited infinitely often (e.g., via exploring starts), and that Assumption 7 holds. Then, the Q-bounds converge to $Q^{*}$ in finite time with probability 1, and the $\pi^{\phi}$ converges to the optimal policy in finite time.
@@
Consider any state-action pair $(x,u)$. The exploring policy ensures that any state-action pair with uncertain Q-bounds is eventually sampled at least once (in finite time). Under Assumption 7, we have:
@@
The reward function set $\mathcal{G}$ is reduced to the true reward function $g$ in finite time, since we have finite state and action spaces. The model uncertainty is thus completely eliminated. By Theorem 6, the exploring policy therefore converges to the optimal policy in finite time with probability 1. ∎
(a) Frozen Lake environment
(b) Cartpole environment
(c) Taxi environment
Figure 2: Evaluation performance of BUMEX versus standard $\epsilon$-greedy Q-learning, Q-learning with UCB exploration, UCRL2, and PSRL across three benchmark environments. Plotted are the median, and the 25th and 75th percentiles of 100 Monte Carlo runs, where the performance is evaluated every 25 training episodes (frozen lake) or 100 training episodes (cartpole, taxi) by averaging the returns obtained by applying the greedy policy for 100 episodes.
Although this paper aims to present a general theoretical framework that can be cast into both finite and infinite state/action spaces, formulating a practical algorithm for the infinite-state, infinite-action case is considerably more challenging. In practice, one must resort to function approximation techniques to parameterize the Q-function. Such a parameterization typically leads to difficulties when performing optimistic max-max optimization because the overestimation inherent in the optimistic operator can result in divergence issues or unstable learning dynamics [16]. The development of such an algorithm is left for future work.
A practical workaround is to construct a finite abstraction of the original system. By discretizing the state and action spaces, one can apply the proposed exploration strategy and Q-bound updates in a tractable manner, with approximation accuracy improving as the discretization becomes finer. This approach faces the same computational scaling challenges as other state-of-the-art finite MDP RL methods such as the ones used in the comparison in the next section.
@@ V RESULTS
We apply the proposed practical algorithm to three benchmark reinforcement learning environments: Frozen Lake, Cartpole, and Taxi from the OpenAI Gym [17].
@@ V-A Frozen Lake Environment
In the Frozen Lake environment, the agent navigates a gridworld to avoid obstacles and reach a goal state, receiving a reward of 1 upon success and 0 otherwise. The transitions are stochastic due to the slippery nature of the ice, with the agent moving in either the intended direction or one of the two perpendicular directions, each with probability $1/3$.
We define the model set to include knowledge about adjacency in the gridworld. In particular, we assume that all transitions for which the true model has probability zero are known, while all others have probabilities between 0 and 1. The reward function set $\mathcal{G}$ contains only the true reward function $g$, assuming full knowledge of the rewards. This model set definition is similar to what a human player might assume when playing the game, i.e., the agent knows the layout of the gridworld and the locations of obstacles and goals, but not the exact transition probabilities. The results are shown in Fig. 2(a).
We compare standard (randomly exploring) $\epsilon$-greedy Q-learning, Q-learning with UCB1 exploration [5], UCRL2 [6], and PSRL [7] to two variants of BUMEX: one without regularization ($L=\infty$) and one with regularization ($L=100$). All experiments use $\xi=1$, $\zeta=0$, $\beta$ as in (14), $c=5$, $\delta=0.05$, $\alpha=0.05$, and $\gamma=0.95$, with $\epsilon$ decaying exponentially from 1 to 0.01. The results indicate that BUMEX converges to the optimal policy quickly and consistently, even when the Q-bound iterations are only performed at the beginning. When performing Q-bound iterations (with regularization) every 100 episodes, BUMEX converges even faster, although this comes at an additional computational cost induced by the repeated calculation of Bellman operators. Note that methods such as UCRL2 and PSRL also rely on repeated Bellman operations with comparable computational costs.
@@ V-B Cartpole Environment
In the Cartpole environment, the agent is tasked with balancing a pole on a cart by moving the cart left or right, receiving a reward of 1 per time step the pole is balanced. The transitions are deterministic and governed by the physics of the cartpole system. We create finite state and action spaces by discretizing the continuous state and action spaces, which introduces stochasticity into the transition model. Furthermore, we perform quadratic reward shaping during the training phase to speed up learning.
@@
We define the model set to include knowledge about the cartpole dynamics, i.e., the cartpole is governed by Newtonian physics. In particular, we assume that the dynamics equations are known up to an unknown mass parameter. We obtain the set $\mathcal{P}$ by looping over the set of possible mass values, obtaining the transition probability tensor for each, then taking the max and min over these tensors to obtain $\underaccent{\bar}{M}$ and $\bar{M}$ in (21). The reward function set $\mathcal{G}$ contains only the true reward function $g$, i.e., we assume full knowledge of the rewards. The results are shown in Fig. 2(b).
@@
We compare the algorithms $\epsilon$-greedy, UCB1, PSRL, BUMEX with $L=\infty$ and $L=200$, with the same $\xi,\zeta,\beta$, $c$, and $\delta$ as in Section V-A (UCRL2 is excluded due to poor performance). We use $\alpha=0.03$, $\gamma=0.97$, and $\epsilon$ decaying exponentially from 1 to 0.001. The results are similar to those in the Frozen Lake environment, showing that the regularized method converges fastest and most consistently, followed by the non-regularized version. The more limited performance gain from regularization in this environment can be attributed to the larger state space dimension.
V-C Taxi Environment
The Taxi environment consists of a more complicated gridworld, in which the agent is tasked with picking up and dropping off passengers at dedicated locations. The agent is rewarded -1 at every timestep, -10 for illegal passenger pickups/dropoffs, and +20 for a successful passenger dropoff. The gridworld contains walls which the agent cannot move through. Note that this environment is deterministic.
Our prior model knowledge largely takes the same form as in Section V-A, i.e., we assume that the agent knows the layout of the gridworld and the locations of obstacles/goals. We assume the agent does not have knowledge of wall placements or the reward function structure.
We compare the same algorithms ($\epsilon$-greedy, UCB1, UCRL2, PSRL, BUMEX with $L=\infty$ and $L=100$) as in the previous sections, with the same $\xi,\zeta,\beta$, $c$, and $\delta$. We use $\alpha=0.03$, $\gamma=0.98$, and $\epsilon$ decaying exponentially from 1 to 0.01. The results are shown in Fig. 2(c) and are consistent with the other two environments. This experiment specifically highlights the benefits of having accurate dynamics knowledge while learning unknown rewards. The proposed method converges very quickly once the deterministic reward function has been learned.
VI CONCLUSIONS AND FUTURE WORK

### v3

@@ III-B1 Choice of Weighting Function $\beta$
$\beta(x,u)=$ $~\text{Prob}[I(x,u)>0]$ (18)
$=$ $~\frac{\max\left(0,\bar{Q}(x,u)-\underaccent{\bar}{V}(x)\right)}{\bar{Q}(x,u)-\underaccent{\bar}{Q}(x,u)}.$
@@
$\beta(x,u)=$ $~\mathbb{E}[I(x,u)]$ (19)
$=$ $~\frac{\left(\max\left(0,\bar{Q}(x,u)-\underaccent{\bar}{V}(x)\right)\right)^{2}}{2\left(\bar{Q}(x,u)-\underaccent{\bar}{Q}(x,u)\right)}.$
@@ III-C Regularization with Observed Data
We propose that by regularizing the optimization over the model set using observed data $\mathcal{D}$, that the bounds converge to the optimal Q-function under mild conditions. We can re-define the Q-bound Bellman updates (5) as
@@
The variable $\lambda$ determines the trade-off between the model set and the observed data. It should increase with the size of the dataset $\mathcal{D}$ to ensure that the model set is close to the observed data, such that
@@
where $T(x,u,x^{\prime})$ the number of observed transitions from state $x$ under action $u$ to state $x^{\prime}$ in the dataset $\mathcal{D}$, and $T(x,u):=\sum_{x^{\prime}}T(x,u,x^{\prime})$.
Additionally, since the rewards are deterministic, the set of reward functions $\mathcal{G}$ can be reduced to a single reward function at a particular state-action pair whenever a reward is observed,
$\mathcal{G}(x_{k},u_{k})=\{r_{k}\}.$ (23)
Together with the regularized model set optimization, we obtain a convergence result which is formalized in the following proposition.
@@
Let $\lambda:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{\geq 0}$ be a regularization parameter satisfying (22), and let $d(P,P_{E})$ be a distance metric such that $d(P,P_{E})\geq 0$ and $d(P,P_{E})=0$ if and only if $P=P_{E}$. Let $\mathcal{P}$ and $\mathcal{G}$ be the set of transition kernels and the set of reward functions, respectively. Assume that the reward function is deterministic, i.e., we can shrink $\mathcal{G}$ through (23) upon observing rewards. If every state-action pair is visited infinitely often, then,
@@
Since every state-action pair is visited infinitely often, $\lambda(x,u)\to\infty$ (by (22)) for every $(x,u)$. The nonnegative metric $d(P,P_{E})$ then forces the regularized optimization in (21) to select $P=P_{E}$, since the regularization term completely dominates the objective. A single visit of every state-action pair is sufficient to ensure that the reward function set $\mathcal{G}=\{g\}$ for every $(x,u)$. By the law of large numbers, $P_{E}$ converges to the true transition kernel $P$, so the regularized update reduces to the standard Bellman update with the true model $(P,g)$. As the standard Bellman operator is a contraction (see, e.g., [1]), it follows that both the lower and upper Q-function bounds converge to $Q^{*}$.
This completes the proof. ∎
@@
Algorithm 1 Q-learning with exploring policy and regularized model set optimization
@@
   Saturate the Q-function $Q$ to the bounds $\underaccent{\bar}{Q},\bar{Q}$.
@@
The regularized model set optimization is incorporated into our proposed method, which is summarized in Algorithm 1, where the data-regularized Q-bound iterations are performed every $L$ episodes. To improve learning efficiency, we employ an $\epsilon$-greedy type strategy, balancing exploration/exploitation. Additionally, we saturate the Q-function to the bounds whenever they are updated. By combining our previous results, we obtain the following convergence guarantee for the exploring policy.
@@
Let $\pi^{\phi}$ be the exploring policy defined in (15) for which we additionally assume that $\zeta$ is positive. Let $\lambda$ satisfy (22). Suppose that $d(P,P_{E})$ is a distance metric with $d(P,P_{E})\geq 0$ and $d(P,P_{E})=0$ if and only if $P=P_{E}$, and that $\mathcal{P}$ and $\mathcal{G}$ are the model sets for the transition kernel and reward function, respectively. Assume that the reward function is deterministic, i.e., we can shrink $\mathcal{G}$ through (23) upon observing rewards, and that every state is visited infinitely often. Then, under Algorithm 1, the Q-bounds converge to the optimal Q-function and the exploring policy $\pi^{\phi}$ converges to the optimal policy with probability 1 as $k\to\infty$.
@@
Almost all the assumptions of Proposition 5 are satisfied. However, the exploring policy $\pi^{\phi}$ in (15) prunes some of the actions, resulting in not every state-action pair being visited infinitely many times. In particular, $\phi$ assigns a positive weight to all actions except actions that are certain and guaranteed to be suboptimal (since $\zeta>0$). Still, since the Q-values of such actions are already certain, this implies that the model set is certain at these state-action pairs. Hence, the exploring policy does not need to explore these actions further, and the regularized model set optimization has the same result regardless of the value of $\lambda$.
For all other state-action pairs, the exploring policy ensures they are explored infinitely often, resulting in $\lambda(x,u)\to\infty$. The regularized model set optimization therefore converges to the true transition kernel for every state-action pair. We obtain (24), similar to Proposition 5. We satisfy (7), which means that the results of Proposition 3 apply. The exploring policy $\pi^{\phi}$ then selects only those actions that are guaranteed to be optimal, implying that it converges to the optimal policy with probability 1.
This completes the proof. ∎
In the next section, we will discuss a practical algorithm derived from the general framework that relates to bounded-parameter MDP models.
IV PRACTICAL ALGORITHM
@@
for known, well-defined lower and upper bounds $\underaccent{\bar}{M},\bar{M}\in[0,1]^{|\mathcal{X}|\times|\mathcal{U}|\times|\mathcal{X}|}$. The reward function set that the true reward function $g$ belongs to is defined as
$\mathcal{G}=\left\{g~\big|~\underaccent{\bar}{g}\leq g\leq\bar{g},\quad\forall~x,u\right\}.$ (27)
In other words, we assume that the true transition kernel and reward function are bounded by known upper and lower bounds.
The distance metric $d$ is chosen to be the Kullback-Leibler divergence between the transition kernel and the observed data, given by
@@
where $M_{E}$ is the empirical transition probability tensor from the observed data $\mathcal{D}$ through $M_{E}(x,u,x^{\prime})=\frac{T(x,u,x^{\prime})}{T(x,u)}$.
@@
We propose to select $\lambda=0$ in the absence of data ($(T(x,u)=0$) and otherwise as a function of system properties and the number of observed transitions from every state-action pair, such that
@@
The scenario presented in this section is a straightforward extension of the tabular Q-learning algorithm, where we can use the Q-function bounds to guide the agent’s exploration. Note that all the convergence properties of Q-learning and those discussed in the previous sections are retained. In some particular cases, we can even guarantee finite-time convergence of the exploring policy to the optimal policy, as shown in the next section.
@@ IV-A Finite-Time Convergence
Furthermore, $\lambda(x,u)$ is chosen such that it completely relies on the data for the deterministic case, that is, $\lambda(x,u)=\infty$ if 1) holds and $T(x,u)>0$.
@@
Let $\pi^{\phi}$ be the exploring policy defined in (15) and let $Q$ be the Q-function from Algorithm 1, using model set $\hat{\mathcal{M}}$ and distance metric $d$ with $d(P,P_{E})\geq 0$ and $d(P,P_{E})=0$ if and only if $P=P_{E}$. Assume that the reward function is deterministic, i.e., we can shrink $\mathcal{G}$ through (23) upon observing rewards. Assume that every state is visited infinitely often (e.g., via exploring starts), and that Assumption 7 holds. Then, the Q-bounds converge to $Q^{*}$ in finite time with probability 1, and consequently the exploring policy $\pi^{\phi}$ converges to the optimal policy in finite time as well.
@@
Consider any state-action pair $(x,u)$. The exploring policy ensures that any state-action pair with uncertain Q-bounds is eventually sampled at least once (in finite time). Under Assumption 7:
@@
Thus, in both cases, the optimistic and pessimistic regularized model set optimization converges to the true transition kernel in finite time. Similarly, the reward function set $\mathcal{G}$ is reduced to the true reward function $g$ in finite time, since we have finite state and action spaces. The Q-bound iterations therefore become equivalent to standard Q-learning updates, since the model uncertainty is completely resolved in finite time. The Q-bounds converge to the optimal Q-function in finite time with probability 1, and using logic similar to the proof of Theorem 6, the exploring policy therefore converges to the optimal policy in finite time with probability 1.
This completes the proof. ∎
Although the general theoretical framework developed in this work applies to both finite and infinite state/action spaces, formulating a practical algorithm for the infinite-state, infinite-action case is considerably more challenging. In practice, one must resort to function approximation techniques to parameterize the Q-function. Such a parametrization typically leads to difficulties when performing optimistic max-max optimization because the overestimation inherent in the optimistic operator can result in divergence issues or unstable learning dynamics [21]. The development of such an algorithm is left for future work.
An effective workaround is to construct a finite abstraction of the original system. By discretizing the state and action spaces, one can apply the proposed exploration strategy and Q-bound updates in a tractable manner, while still capturing the essential dynamics of the underlying continuous problem.
@@ V RESULTS
In this section, we demonstrate the merits of the proposed practical algorithm by applying it to two benchmark reinforcement learning environments, namely the Frozen Lake and Cartpole environments from the OpenAI Gym [2].
@@ V-A Frozen Lake Environment
In the Frozen Lake environment, the agent navigates a gridworld to avoid obstacles and reach a goal state, receiving a reward of 1 upon success and 0 otherwise. The transitions are stochastic due to the slippery nature of the ice, with the agent moving in each perpendicular direction with probability $1/3$.
We define the model set to include knowledge about adjacency in the gridworld. In particular, we assume that all transitions for which the true model has probability zero are known, while all others have probabilities between 0 and 1. The reward function set $\mathcal{G}$ contains only the true reward function $g$, assuming full knowledge of the rewards. This model set definition is similar to what a human player might assume when playing the game, i.e., the agent knows the layout of the gridworld and the locations of obstacles/goals, but not the exact transition probabilities. The results are shown in Fig. 2.
Figure 2: Evaluation performance of the proposed exploration strategy versus standard $\epsilon$-greedy Q-learning in the Frozen Lake environment. Plotted are the median, and the 5th and 95th percentiles of 100 Monte Carlo runs, where the performance is evaluated every 25 episodes by averaging 100 runs of the greedy policy.
We compare standard (randomly exploring) $\epsilon$-greedy Q-learning to two variants of our algorithm: one using the nonregularized exploring policy ($L=\infty$) and one using the regularized version with $L=50$. All experiments use $\xi=1$, $\zeta=0$, $\beta$ as in (19), $c=5$, $\delta=0.05$, $\alpha=0.05$, and $\gamma=0.95$, with $\epsilon$ decaying exponentially from 1 to 0.01. The results indicate that the proposed method converges to the optimal policy much more quickly and consistently, even when the Q-bound iterations are only performed at the beginning. When performing Q-bound iterations every 50 episodes, the proposed method converges even faster, although this comes at an additional computational cost induced by the repeated calculation of Bellman operators.
@@ V-B Cartpole Environment
In the Cartpole environment, the agent is tasked with balancing a pole on a cart by moving the cart left or right, receiving a reward of 1 per time step the pole is balanced. The transitions are deterministic and governed by the physics of the cartpole system. We create finite state and action spaces by discretizing the continuous state and action spaces, resulting in a stochastic transition model. Furthermore, we perform quadratic reward shaping during the training phase to speed up learning.
@@
We define the model set to include knowledge about the cartpole dynamics, i.e., the cartpole is governed by Newtonian physics. In particular, we assume that the dynamics equations are known up to an unknown mass parameter. We obtain the set $\mathcal{P}$ by looping over the set of possible mass values, obtaining the transition probability tensor for each, then taking the max and min over these tensors to obtain $\underaccent{\bar}{M}$ and $\bar{M}$ in (26). The reward function set $\mathcal{G}$ contains only the true reward function $g$, i.e., we assume full knowledge of the rewards. The results are shown in Fig. 3.
Figure 3: Evaluation performance of the proposed exploration strategy versus standard $\epsilon$-greedy Q-learning in the Cartpole environment. Plotted are the median, and the 5th and 95th percentiles of 100 Monte Carlo runs, where the performance is evaluated every 100 episodes by averaging 100 runs of the greedy policy.
@@
We compare the same algorithms ($\epsilon$-greedy, $L=\infty,L=500$) as in Section V-A, with the same $\xi,\zeta,\beta$, and $\delta$. We use $c=100,\alpha=0.03$, $\gamma=0.97$, and $\epsilon$ decaying exponentially from 1 to 0.001. The results are similar to those in the Frozen Lake environment, showing that the regularized method converges fastest and most consistently, followed by the non-regularized version, and then the standard $\epsilon$-greedy Q-learning algorithm. The performance gain due from the regularization is more limited in this environment, which can be attributed to the larger state space dimension.
VI CONCLUSIONS AND FUTURE WORKS